OpenStudy (anonymous):

rational expression equation help: (x+4)/(x-5) = -14/13

OpenStudy (jdoe0001):

just cross multiply and solve for "x" :)

OpenStudy (anonymous):

can you show steps? when I cross multiplied I didn't get the correct answer

OpenStudy (jdoe0001):

well, after you cross multiply you have to "isolate" x

OpenStudy (anonymous):

i got -14x + 136 = -13x - 87

OpenStudy (jdoe0001):

hmmm, well 13 * (x+4 ) = 13x + 52 - 14 ( x-5) = -14x + 70

OpenStudy (anonymous):

whoops numbers from another problem.

OpenStudy (anonymous):

why isn't the 13 negative as well as the 14?

OpenStudy (jdoe0001):

when you have a negative fraction, either the numerator or the denominator is negative only \(\bf \large {\frac{a}{b} \implies\frac{a}{b}\\ \frac{-a}{-b} \implies\frac{a}{b}\\ \frac{-a}{b} \implies-\frac{a}{b}\\ \frac{a}{-b} \implies -\frac{a}{b} } \)

OpenStudy (jdoe0001):

in my cases I used the -14, instead of the -13 if I had used -13, then I'd have needed to use +14

OpenStudy (anonymous):

The actual answer is 2/3. I got -27x=-20. so 20/27. that isn't 2/3

OpenStudy (jdoe0001):

13x + 52 = -14x + 70 # adding 14x to both sides 27x +52 = 70 # substracting 52 from both sides 27x = 18 thus x = 18/27 # simplifying by dividing each by 9 x = 2/3

OpenStudy (anonymous):

Argh. missed the 52 and wrote 50. thanks for your help.

OpenStudy (jdoe0001):

yw